Status Report

NASA Space Station On-Orbit Status 31 August 2004

By SpaceRef Editor
August 31, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 31 August 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below. 

Working mostly in the DC1 “Pirs” module, the crew continued Orlan suit preparations for EVA-11 on Friday, focusing first on valve functionality verification on the suits and their BSS interface control units.  Most of the operations involving Orlan replaceable equipment checkout & installation on today’s schedule had already been completed yesterday.    [Orlan consumable ORUs in the backpacks include: primary BK-3 oxygen (O2) tank, 825M3 batteries, LP-9 lithium hydroxide (LiOH) canisters for carbon dioxide (CO2) absorption, moisture collectors, FOR feedwater line filters, BOS filtration & separation unit filters and IK CO2 measurement units with filters.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Also completed by the crew yesterday were final fit checks of their spacesuits (Padalka: #25, with red markings and BRTA radio telemetry unit #13; Fincke: #26 with blue markings and BRTA #18).   [The crew used these suits during EVA-10.  By means of straps and cables, they can be adjusted for size in body length, lateral size, hip, calf, arm length and shoulder dimension.]

U.S. gear installed in the Orlans includes DIDBs (disposable in-suit drink bags) and EHIP (EMU helmet interchangeable portable) lights, borrowed from U.S. EMU/spacesuits and jerry-rigged on the Orlans.  For powering the EHIPs, Mike Fincke today recharged four NiMH (nickel metal hydride) batteries (#1009, #1016, #1018 & #1020).

Gennady and Mike also checked out their PKO-BETA08 ECG (electrocardiogram) harnesses on the GAMMA-1 biomedical examination panel (PKO) before they set up for a thorough test of Orlan BRTA telemetry & communications from the DC1.  Testing of communication links included VHF/voice and biomedical electrode belt and telemetry hookups via the BSS (later by the wireless in-suit Tranzit-B radio telemetry system) for vital signs and equipment monitoring.   [The DC1 comm control panel (InPU) was reconfigured to command VHF off, in order to prevent extraneous radio noise outside RGS (Russian ground site) passes from entering the system while working in the Orlan suit.  Per MCC-H request, the RSA2 audio link was patched into the U.S. voice channel “Public Call 2” to provide EVA comm from the USOS.]

CDR Padalka retrieved a spare BNP portable air repress bottle from the “divan” of the Soyuz orbital module and installed it with Velcro (Russian: Aramid) tape in the DC1’s instrumentation “corridor”.   [This additional BNP is used as standard protection against the contingency of the KVD repress valve to the corridor being stuck Closed during the post-ingress repressurization of the airlock.]

FE/SO Fincke worked on the HRF (human research facility) rack, checking out the MedOps cardiac defibrillator.   [This periodic routine task is scheduled as soon as possible from Expedition start and every 60 days thereafter.  For the checkout, the defib is connected to the 120V outlet, equipped with its battery and then allowed to charge, for about five seconds, to a preset energy level (e.g., 100 joules).  After the button-triggered discharge, a console indicator signals success or failure of the test.  The pacing signal is to be downlinked via S-band for 2 min.  The HRF is powered down subsequently.]

Also on the HRF PC, the FE performed standard folder/file management for the ADUM (Advanced Diagnostic Ultrasound in Microgravity), moving accumulated scanning files between hard disk partitions by copying them from the filled-up D drive to C drive.

After EXPRESS rack 2 (ER2) and ER5 were powered up by the ground, Fincke activated the ER5 laptop computer (ER5 ELC).  He then initiated a 3-hr. session of the Serial Network Flow Monitor (SNFM) application for the ELC to capture packet data traffic on the LAN-2 science Ethernet network during SAMS (space acceleration measurement system) rack-to-rack comm traffic, for analysis on the ground.

Mike also completed a 20-min. task to cover redundant fire ports in the Node and Lab with gray tape to reduce the possibility of crew confusion during an on-orbit fire event.

Gennady performed the routine inspection of the SM’s SOZh environment control & life support system, which today included the weekly servicing of the BRPK-1 air/water condensate separator, and later prepared the regular IMS (inventory management system) “delta” file for export/import to the IMS databases.

The crew completed its daily physical exercise on TVIS treadmill, RED resistive exerciser, and VELO stationary bike with load trainer.

Mike Fincke was provided with a new list of options for his next “Saturday Science” program for his choice.   [Suggested by POIC (Payload Operations & Integration Center) are an ISSI (In-Space Soldering Investigation) test (#3 of 5), an MFMG (Miscible Fluids in Microgravity) test, operations with the HEAT experiment (investigating heat transfer performance of heat pipes), and another CFE (Capillary Flow Experiment) run.]

At ~2:25pm, lasting for several hours, TsUP/Moscow plans to perform a regular standard calibration session on the Soyuz TMA-4/8S vehicle’s critical BILU linear accelerometer unit via BITS2-12 onboard telemetry system in VD-SU control system mode, using RGS for command uplink and U.S. S-band for telemetry downlink.

Yesterday, Fincke successfully removed and replaced the service life-expired CMOS (complementary metal-oxide semiconductor) battery of SSC1 (station support computer 1), the first time for this procedure to be attempted on orbit.   SSC1 can now be considered a backup laptop for a PCS (portable computer system) machine if needed.  CMOS battery R&R for the SSC Router (shell #6007), another IBM ThinkPad 760XD, is scheduled for tomorrow.

Early on 8/30 (yesterday), at 12:47am EDT, the Elektron experienced another bubble-caused shutdown.  TsUP linked up a new procedure for the crew to execute for restarting Elektron, which they accomplished successfully at 4:16am in 16A mode.  Later, performance level was raised to 20A, as it was prior to the shutdown.

Major upcoming events:

  • EVA-11 — 9/3 (hatch open 12:50pm EDT);
  • Soyuz 9S launch — 10/9;
  • Soyuz 9S dock — 10/11;
  • Soyuz 8S undock/land — 10/19;
  • Soyuz 9S relocate — 11/18;
  • Progress 16P launch — 11/24.

Expedition 9 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Today’s CEO photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were Hurricane Frances, western Atlantic (Dynamic Event.  Frances is currently a Category 3 hurricane, and is predicted to head towards the eastern Florida coast.  The eye should have been located approximately 2 degrees to the left of track), Internal waves, Tuamotu Archipelago (the ISS overpass provided an opportunity to capture internal waves in the western archipelago.  Some clouds were present but weather should have been relatively clear.  The sunglint point was to the left of track), and Tuamotu-Austral Islands (the overflight provided opportunities for both the northwestern (right of track and behind ISS) and the southeastern (parallel to and left of track) islands.  Due to the current XPOP orientation the northwestern islands may have been easier to photograph).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

U.S. & Russian Segment Status  (as of today, 12:50pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On. SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is still considered failed).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 738; temperature (deg C) — 26.8; ppO2 (mmHg) — 161.0; ppCO2 (mmHg) — 3.6.
  • SM Transfer Compartment:  Pressure (mmHg) — 736; temperature (deg C) — 20.9.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 21.7.
  • Node:  Pressure (mmHg) — 738.2; temperature (deg C) — 24.1 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 739.8; temperature (deg C) — 24.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 740.0; temperature (deg C) — 25.5; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, “sun slicer”, i.e., drag reduction-biased by 47 deg angle (2B: +47, 4B: -47).
  • SM batteries:  All batteries (8) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #1 is off line; battery #2 is in “Cycle” mode.  All other batteries (4) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems (C&DH)

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System (PS):

  • Total propellant load available: 4239 kg (9345 lb) as of 8/26/04;  [SM(552) + FGB(3407) + Progress M(280)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems (ACS):

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 180.5 deg, pitch: -6.9 deg., roll: 0 deg]), with CMG TA (thruster assist) until 9/2 for EVA-11, then back to XPOP until next reboost (9/22).

Communications & Tracking Systems (C&T):

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 7:23am EDT [= epoch]):

  • Mean altitude — 361.3 km
  • Apogee height — 365.1 km
  • Perigee height — 357.4 km
  • Period — 91.77 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0005707
  • Solar Beta Angle — 34.2 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss in last 24 hours — 77 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 33014

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.